# An Survey of the Evolution of Language Model-Based Dialogue Systems

### v1

@@ Abstract
Dialogue systems (DS), including the task-oriented dialogue system (TOD) and the open-domain dialogue system (ODD), have always been a fundamental task in natural language processing (NLP), allowing various applications in practice. Owing to sophisticated training and well-designed model architecture, language models (LM) are usually adopted as the necessary backbone to build the dialogue system. Consequently, every breakthrough in LM brings about a shift in learning paradigm and research attention within dialogue system, especially the appearance of pre-trained language models (PLMs) and large language models (LLMs). In this paper, we take a deep look at the history of the dialogue system, especially its special relationship with the advancements of language models. Specifically, we delineates the evolutionary trajectory of LM-based dialogue systems by categorizing it into four distinctive stages, each marked by pivotal breakthroughs in language models: 1) early stage, which witnesses predominantly statistical LMs, resulting in rule-based or machine-learning-driven DS; 2) the independent development of TOD and ODD based on neural language models (NLM; such as LSTM and GRU), since NLMs lack intrinsic knowledge in their parameters; 3) fusion between different types of dialogue systems with the advert of pre-trained language models (PLMs), starting from the fusion between four sub-tasks within TOD, and then TOD with ODD; and 4) current LLM-based dialogue system, wherein LLMs can be used to conduct TOD and ODD seamlessly. In this way, our survey offers a systematic perspective, categorizing the stages in a chronological order aligned with LM breakthroughs, providing a comprehensive review of state-of-the-art research outcomes. What’s more, we turn our attention to emerging topics and engage in a discussion on open challenges, providing valuable insights into the future directions for LLM-based dialogue systems. In summary, this survey delves into the dynamic interplay between language models and dialogue systems, unraveling the evolutionary path of this essential relationship. Through this exploration, we pave the way for a deeper comprehension of the field, guiding future developments in LM-based dialogue systems.
Index Terms:
@@
Task-oriented Dialogue System, Open-domain Dialogue System, Pre-trained Language Models, Large Language Models.
1 Introduction
@@
Building a conversational intelligent system has always been a fundamental objective in the realm of natural language processing [1, 2]. The ability to interact naturally and seamlessly with machines has opened up new avenues for human-machine communication, leading to transformative applications across different industries, from Siri, Xiaoice [3] to the latest New Bing  [^1] https://www.bing.com/new and Google Bard  [^2] https://bard.google.com/. At the heart of this evolution is the pivotal role that dialogue systems play, which are designed to facilitate interactions characterized by harmlessness [4], helpfulness [5], trustworthiness [6], and personalization [7]. These systems aim to emulate human-to-human conversations, thereby offering the potential to enhance user experiences, streamline tasks, and provide personalized assistance across domains such as customer support, virtual assistants, healthcare, education, and more, depending on the types of dialogue context [8] as shown in Table I.
@@
Specifically, there are two major types of dialogue in practice: task-oriented dialogue and chit-chat. Task-oriented dialogue revolves around assisting users in achieving specific tasks or goals, such as making reservations or booking tickets. In this context, the dialogue system acts more like a helpful assistant, providing relevant information and guidance to users [9, 2]. Conversely, chit-chat involves casual and informal conversations, mainly aimed at establishing and maintaining social connections. Here, the dialogue system adopts the role of a friendly chatbot, engaging users in a more relaxed and conversational manner [1]. Based on their distinct roles, two types of dialogue systems have been proposed: Task-oriented Dialogue System (TOD) and Open-domain Dialogue System (ODD). The TOD is designed to efficiently handle task-oriented conversations, guiding users towards accomplishing specific objectives by detecting user intentions, tracking dialogue state, making suitable actions, and responding accordingly. In contrast, the ODD is intended for open-domain interactions, enabling free-flowing conversations on a wide range of topics by directly mapping the dialogue context into the response, without a predefined task or goal. These two types cater to different use cases and user needs, showcasing the versatility and applicability of dialogue systems in diverse scenarios. Thus, most of the previous works (before $\sim$2020, Figure 2) aim to design and build these dialogue systems independently using different backbones considering their different roles [2, 1].
Figure 1: The development of LM-based Dialogue Systems, which can be divided into four distinct stages, each marked by significant breakthroughs that have paved the way for advancements in the field: 1) Early Stage (1966 - 2015); 2) The Independent Development of TOD and ODD (2015 - 2019); 3) Fusions of Dialogue Systems (2019 - 2022); 4) LLM-based DS (2022 - Now).
@@
However, the revolutionary progress in Language Models (LMs) has sparked a transformative evolution in dialogue systems, reshaping their very foundation. The core objective of LMs is to predict word sequence probabilities has brought about a profound impact on dialogue systems [10, 11], empowering them to learn world knowledge from the pre-training corpus [12] and generate more contextually relevant and helpful responses [13]. Concurrently, each milestone in LM development has acted as a catalyst for transformative changes in dialogue systems, playing a pivotal role in their rapid evolution and ongoing trends, especially the appearance of latest large language models (LLMs), such as ChatGPT. Rather than dwelling on the development of language models, our focus lies in the enthralling journey of dialogue systems undergoing this transformative process, which can be divided into four major development stages corresponding to the four stages of LM [14] as shown in Figure 1:
@@
TABLE I: Different types of dialogues in TOD and ODD. We keep the last turn of users for conciseness and brevity.

### v2

# A Survey of the Evolution of Language Model-Based Dialogue Systems: Data, Task and Models

@@ Abstract.
Dialogue systems (DS), including the task-oriented dialogue system (TOD) and the open-domain dialogue system (ODD), have always been a fundamental task in natural language processing (NLP), allowing various applications in practice. Owing to sophisticated training and well-designed model architecture, language models (LM) are usually adopted as the necessary backbone to build the dialogue system. Consequently, every breakthrough in LM brings about a shift in learning paradigm and research attention within dialogue system, especially the appearance of pre-trained language models (PLMs) and large language models (LLMs). In this paper, we take a deep look at the history of the dialogue system, especially its special relationship with the advancements of language models. Thus, our survey offers a systematic perspective, categorizing different stages in a chronological order aligned with LM breakthroughs, providing a comprehensive review of state-of-the-art research outcomes. What’s more, we turn our attention to emerging topics and engage in a discussion on open challenges, providing valuable insights into the future directions for LLM-based dialogue systems. In summary, this survey delves into the dynamic interplay between language models and dialogue systems, unraveling the evolutionary path of this essential relationship. Through this exploration, we pave the way for a deeper comprehension of the field, guiding future developments in LM-based dialogue systems.
@@
Task-oriented Dialogue System, Open-domain Dialogue System, Pre-trained Language Models, Retrieval-augmented Generation
This project is partially supported by HKSAR ITF Project ITT/008/22LP
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Computing methodologies Natural language generation
1. Introduction
@@
Building a conversational intelligent system has always been a fundamental objective in the realm of natural language processing (Huang and et al., 2020; Zhang and et al., 2020b). The ability to interact naturally and seamlessly with machines has opened up new avenues for human-machine communication, leading to transformative applications across different industries, from Siri, Xiaoice (Zhou et al., 2020) to the latest New Bing  [^1] https://www.bing.com/new and Google Bard  [^2] https://bard.google.com/. At the heart of this evolution is the pivotal role that dialogue systems play, which are designed to facilitate interactions characterized by harmlessness (Ouyang et al., 2022), helpfulness (Wang and et al., 2023b), trustworthiness (Huang et al., 2021), and personalization (Salemi et al., 2023). These systems aim to emulate human-to-human conversations, thereby offering the potential to enhance user experiences, streamline tasks, and provide personalized assistance across domains such as customer support, virtual assistants, healthcare, education, and more, depending on the types of dialogue context (Liu et al., 2022b). Throughout the entire evolution of conversational intelligent systems, two factors have been pivotal: data and model.
@@
Data. Specifically, there are two major types of dialogue data in practice: task-oriented dialogue and chit-chat. Task-oriented dialogue revolves around assisting users in achieving specific tasks or goals, such as making reservations or booking tickets. In this context, the dialogue system acts more like a helpful assistant, providing relevant information and guidance to users (Hosseini-Asl et al., 2020; Zhang and et al., 2020b). Conversely, chit-chat involves casual and informal conversations, mainly aimed at establishing and maintaining social connections. Here, the dialogue system adopts the role of a friendly chatbot, engaging users in a more relaxed and conversational manner (Huang and et al., 2020). Based on their distinct roles, two types of dialogue systems have been proposed: Task-oriented Dialogue System (TOD) and Open-domain Dialogue System (ODD). The TOD is designed to efficiently handle task-oriented conversations, guiding users towards accomplishing specific objectives by detecting user intentions (natural language understanding, NLU), tracking dialogue state (dialogue state tracking, DST), making suitable actions (dialogue policy learning, DPL), and responding accordingly (natural language generation, NLG). In contrast, the ODD is intended for open-domain interactions, enabling free-flowing conversations on a wide range of topics by directly mapping the dialogue context into the response, without a predefined task or goal. These two types cater to different use cases and user needs, showcasing the versatility and applicability of dialogue systems in diverse scenarios. Thus, most of the previous works (before $\sim$2020) aim to design and build these dialogue systems independently using different backbones considering the intrinsic characteristics of different types of data (Zhang and et al., 2020b; Huang and et al., 2020).
@@
Model. The revolutionary advancements in language models (LMs) have catalyzed a transformative evolution in numerous natural language tasks, including dialogue systems, reshaping their very foundation. The core objective of LMs is to predict word sequence probabilities has brought about a profound impact on dialogue systems (Tran and et al., 2017; Radford et al., 2018), empowering them to learn world knowledge from the pre-training corpus (Petroni et al., 2019) and generate more contextually relevant and helpful responses (Zhang et al., 2020b). Concurrently, each milestone in LM development brings paradigm shift in dialogue systems, playing a pivotal role in their rapid evolution and ongoing trends, especially the appearance of latest large language models (LLMs), such as ChatGPT.
Rather than focusing solely on the development of dialogue systems from a data perspective or the advancement of language models from a model perspective, as previous surveys have done (Chen and et al., 2017; Ni et al., 2023), we emphasize the dynamic interaction between data and models. This survey highlights the fascinating evolution of dialogue systems through this transformative interplay, showcasing a clear path from traditional rule-based dialogue system towards current dominated LLM-based dialogue systems. This whole progression can be categorized into four major stages, corresponding to the four developmental phases of language models as illustrated in Figure 1.
Figure 1. The development of LM-based Dialogue Systems, which can be divided into four distinct stages, each marked by significant breakthroughs that have paved the way for advancements in the field: 1) Early Stage (1966 - 2015); 2) The Independent Development of TOD and ODD (2015 - 2019); 3) Fusions of Dialogue Systems (2019 - 2022); 4) LLM-based DS (2022 - Now).
Early stage – Statistical language models (SLMs). The first dialogue system — Eliza, was proposed by MIT in 1966, which was even earlier than the rise of SLMs in the 1990s, allowing plausible conversations between humans and machines. Subsequently, a series of dialogue systems were constructed, including a travel agent aiding the client to make a simple return trip to a single city in California (Bobrow et al., 1977) and a rule-based chatbot configured with a persona, employing heuristic pattern matching rules to address inquiries about ages, hobbies, interests, and etc (Wallace, 2009)  [^3] https://web.njit.edu/~ronkowit/eliza.html. In addition to academic endeavors, more and more companies pay attention to the field of dialogue systems and have developed their own products. For example, IBM has developed WATSON, Apple has Siri, and Microsoft has XiaoIce. These early virtual assistants have limited functionalities and follow a rigid process, primarily assisting users with single and simple tasks like scheduling appointments, setting reminders, and providing basic answers to questions. At this stage, most of these dialogue systems are task-oriented, the prevailing nature of these dialogue systems is predominantly task-oriented, employing either modular-based approaches or machine-learning methods based on SLMs.
The independent development of TOD and ODDNeural language models (NLMs). Around 2015, there was a sea change when the sequence to sequence (seq2seq) framework was proposed, in which salient features were learned jointly with the training of the model — NLM (Bengio et al., 2000; Mikolov et al., 2010). Specifically, NLMs characterize the probability of word sequences by neural networks, e.g., two typical recurrent neural networks (RNNs) such as long short-term memory networks (LSTM) (Hochreiter and Schmidhuber, 1997) and gated recurrent units (GRU) (Cho and et al., 2014). These NLMs serve as the backbone to map variable-length input sequences to variable-length output sequences which is achieved by employing two main components: an encoder and a decoder. This paradigm shift facilitated the emergence of Open-domain Dialogue (ODD) systems ($\sim$2015), garnering attention due to a well-defined end-to-end framework enabled by the seq2seq network. Simultaneously, NLMs were also applied to TOD, with the goal of achieving better performance and performing complex tasks in multiple domains (Peng and et al., 2017; Wen et al., 2017b).
Fusions of Dialogue Systemspre-trained language models (PLMs). Traditional NLMs requires extensive hand-crafted features selection and domain-specific knowledge, making them cumbersome and time-consuming to develop for each task. However, with the advent of deep learning and large-scale language corpora, pre-trained language models (PLMs) emerged as a groundbreaking solution. These models are pre-trained on massive amounts of unlabeled text data, capturing rich semantic and syntactic patterns (Devlin and et al., 2019; Radford et al., 2018, 2019). By leveraging pre-training, subsequent fine-tuning on specific tasks becomes more accessible, enabling the models to adapt quickly and achieve impressive results with minimal task-specific data. In the realm of dialogue systems, DialoGPT exemplifies this approach by undergoing pre-training on 147 million conversation-like exchanges extracted from Reddit comment chains and then finetuning on a few training examples to generate more relevant, informative, and context-consistent responses (Zhang et al., 2020b). This paradigm, commonly known as pre-train and finetune (Liu and et al., 2023a), leads to the resulting model as pre-trained dialogue model (PDM) that undergoes fine-tuning using a dialogue corpus. It is worth noting that the major differences between PDMs lies in the the data used for training and the backbone language models adopted, such as PLATO (Bao et al., 2020; Bao and et al., 2021), BlenderBot (Roller et al., 2021), Meena (Adiwardana et al., 2020), Pangu-bot (Mi et al., 2022), PLUG (Bi and et al., 2020), and etc. During this stage, there is an increasing focus on the different level of fusion: 1) different sub-tasks inside TOD, resulting in end-2-end TOD (Qin and et al., 2023), 2) TOD with ODD, resulting in unified dialogue system.
LLM-based dialogue systemLarge language models (LLM). Recognizing the correlation between increased pre-training corpus size and model size with enhanced performance across various NLP tasks (Kaplan et al., 2020), researchers try to scale the model size and the size of pre-training corpus simultaneously to be more sample efficient, enabling the model to learn more intricate patterns and representations from the data. Thus, the PLMs become Large Language Models (LLMs), such as GLM (Du et al., 2022), LLaMA (Touvron et al., 2023a), and InstructGPT (Ouyang et al., 2022). Due to the larger model size and the utilization of extensice, high-quality pre-training corpora, these LLMs offer unprecedented capabilities in many language understanding and generation tasks, including but not limited to question-answering and named entity recognition, thus reshaping the landscape of dialogue systems. With further instruction-tuning using data as (instruction, input, output), these LLMs are capable of answering diverse questions and following various instructions such as writing an email and telling a joke. At this moment, the LLMs can be directly used as the dialogue systems. Moreover, finetuning the LLM with in-domain data, i.e., dialogue/conversational corpus, can further enhance its conversational ability. For example, the evolution of GPT-3.5 to InstructGPT, and then ChatGPT exactly follows this path. Similarly, ChatGLM (Zhang and et al., 2023b) is also finetuned using publicly available Chinese dialogue datasets based on the corresponding LLM – GLM. To enhance alignment with human preferences and values, the incorporation of reinforcement learning becomes imperative, particularly in the final stages of model refinement. This involves leveraging feedback from either human evaluators (Ouyang et al., 2022) or automated systems (Wang et al., 2022) to iteratively improve the model’s performance and ensure its adherence to desired criteria. After all, this type of LLM-based dialogue system (LLM-based DS) performs well at both task-oriented dialogue and chit chat, making it an ideal foundation for a general conversational AI.
It is worth noting that the development of LM-based Dialogue System is a ongoing and evolving process, and the stages are not rigidly separated by specific turning points. Instead, there are overlaps and continuous advancements across these stages, particularly in the later stages. Generally, with the advancement of LM, the boundary between TOD and ODD and the boundary between dialogue model and language model have become increasingly blurred, ushering in a new era of LLM-based dialogue system. Despite that the existing studies have thoroughly examined the development and challenges associated with different types of dialogue systems, such as TOD (Zhang and et al., 2020b) and ODD (Huang and et al., 2020), there still remains a notable gap in the literature when it comes to offering a holistic perspective on the evolutionary trajectory of dialogue systems, particularly in light of the advancements in LM. By critically analyzing and understanding the impact of pivotal milestones, we can gain valuable insights into the possibilities that await and the transformative impact dialogue systems may have on diverse domains and applications in the foreseeable future. This survey aims to present a comprehensive overview of the realm of LM-based dialogue systems  [^4] We have made thorough efforts to cover a wide range of works, but acknowledge the possibility of omissions. Readers are encouraged to provide suggestions for any missed contributions or errors. We are committed to updating the article with new approaches or definitions proposed by the community., shedding light on the prominent directions which the field is currently progressing. This entails a meticulous review of the existing literature, aimed at elucidating the paradigm shifts observed within different stages of LM-based dialogue systems. These shifts encompass a range of integrations across various levels, notably encompassing the data, model and task dimensions. In order to highlight this evolution path, we will not focus on specific minor modifications of model for different tasks (i.e., LSTM and GRU) but instead showcase how these fusion happens with advancement of language models (Tay and et al., 2022). We emphasize this is more important at the era of large language models, since most of previous modifications at the large language model level becomes inefficient and infeasible, resulting in unpredictable consequence.
The remainder of this survey is organized as follows: Section 2 provides the description of different types of dialogues, unified task definition and corresponding commonly used metrics and benchmarks in dialogue system, followed by the early development of two types of LM-based dialogue systems: TOD and ODD in Section 3. Sections  4 review and summarize the evolution of different dialogue systems when PLMs are scaled to LLMs. We then investigate emerging trends and open challenges in LLM-based DS (See Section 5). Then, Section 6 discusses the open problems for LLM-based dialogue systems. Finally, we conclude the survey in Section 7 by summarizing the major findings and discussing the remaining issues for future work.
@@
Table 1. Different types of dialogues in TOD and ODD. We keep the last turn of users for conciseness and brevity.